OpenStudy (anonymous):

Given that f(x) = 6x + 2 and g(x) = the quantity of 2x plus 4 divided by 5 solve for g(f(1)).

OpenStudy (anonymous):

f(x) = 6x + 2 g(x) = (2x + 4)/5 So first find f(1) = 6(1) + 2 = 8 and g(8) = 20/5 = 4.

OpenStudy (anonymous):

thank you so much